# Generative Artificial Intelligence-Supported Pentesting: A Comparison between Claude Opus, GPT-4, and Copilot

### v1

@@ 6 Conclusion and Future Work
In this study, we examined the role of Generative Artificial Intelligence (GenAI) tools in supporting a penetration testing (pentesting) process based on the Penetration Testing Execution Standard (PTES). Specifically, we evaluated the use of Claude Opus, GPT-4, and Copilot across each phase of PTES within a virtualized environment.
@@
In addition to supporting professional pentesters, these tools can also serve as valuable educational aids, providing novice pentesters with examples of attack scenarios, mitigation strategies, and detailed reports for training purposes. To maximize the effectiveness of GenAI tools, future work should explore the integration of these tools with established pentesting frameworks and technologies. By combining the strengths of AI with proven tools, the pentesting process could be significantly enhanced, improving both the speed and accuracy of assessments.

### v2

@@ 6 Conclusion and Future Work
In this study, we examined the role of Generative Artificial Intelligence (GenAI) tools in supporting a penetration testing (pentesting) process based on the Penetration Testing Execution Standard (PTES). Specifically, we evaluated the use of Claude Opus, GPT-4 Turbo, and Copilot across each phase of PTES within a virtualized environment.
@@
In addition to supporting professional pentesters, these tools can also serve as valuable educational aids, providing novice pentesters with examples of attack scenarios, mitigation strategies, and detailed reports for training purposes. To maximize the effectiveness of GenAI tools, future work should explore the integration of these tools with established pentesting frameworks and technologies. By combining the strengths of GenAI with proven tools, the pentesting process could be significantly enhanced, improving both the speed and accuracy of assessments. Another critical factor influencing outcomes is the human skill involved in crafting prompts. The effectiveness of LLMs in pentesting depends significantly on how precisely the pentester formulates queries and structures context. As such, the productivity gains offered by GenAI tools are partially contingent on the pentester’s experience and knowledge in both cybersecurity and prompt engineering. It is also important to point out that using these GenAI tools in real-world pentesting poses legal and ethical risks, such as unsafe command suggestions, data privacy violations, and inaccurate outputs. Pentesters should validate AI-generated commands, avoid sharing sensitive data, maintain human oversight, and document AI usage. Clearer guidelines are needed as the technology evolves.